# C. Röver, Q. Wu, A. Loos, T. Friede

### v1

@@ 6 Discussion
In evidence synthesis and meta-analysis contexts, the one-to-one correspondence of results from separate experiments or studies is often questionable, due to differences in designs, follow-up times, or definition of AEs. In addition, differences in capturing adjudicating and reporting clinical events can increase between-study heterogeneity. The anticipation and consideration of (potential) heterogeneity between data sources hence is commonly advocated.41, 42 On the technical side, implementation of analysis models accounting for heterogeneity usually leads to more cautious or conservative inferences.34, 35 We have extended Holzhauer’s proposed meta-analysis model for synthesizing adverse event counts3 to incorporate a random-effect (RE) component in order to safeguard against overly optimistic or overconfident conclusions. When Holzhauer presented the original model, it was applied in an example case18 that had already triggered some controversy,26, 43, 44 and he was able to advance the discussion by leveraging additional external evidence via an informative MAP prior.4 We were able to demonstrate that the consideration of heterogeneity and use of an RE model also makes a difference here; using the rosiglitazone example as well as an additional application in oncology, we could show the impact of model specification details on the conclusions for these particular applications as well as in some limited Monte Carlo studies. Use of meta-analysis methods as well as consideration of heterogeneity between data sources allows to broaden the evidence base for decision making beyond the limits of single clinical trials that are usually powered with a focus on certain specific efficacy endpoints. Consideration of additional data will allow for more definite statements about harms and benefits, while allowing for exploratory investigations into wider ranges of questions. In the examples, very different disease areas were considered. Depending on the disease, only very specific AEs of interst qualify for such summaries as otherwise the noise from the anticipated events (i.e. those seen in the population of interest regardless of investigational treatment) may be too strong.
@@
Note that model specification in terms of a random “control” response (here: $\lambda_{i0}$) and a random treatment effect ($\varphi_{i}$) implies greater variability in the treatment responses compared to the control arms. In general, this is an often neglected feature, in the oncology example (Section 4.2), this was explicitly exploited in order to allow for greater variability for the (rather heterogeneous) control arms. Alternatively, one could specify a “symmetric” model variation as in Jackson et al. (2018),45 which may be preferrable in some applications. Another common issue in oncology applications is the use of active comparators, so that the extension to network meta-analysis methods may be an option; as long as only two-armed trials are included, this may be accommodated in a rather simple meta-regression model.46 Another aspect in which a more realistic model specification might be worth considering is a possible extension to a competing-risks model.47
@@ Appendix A Heterogeneity posteriors
Figure 8 illustrates the posteriors for the heterogeneity parameters ($\eta$) for two selected analyses of the rosiglitazone and oncology data sets from Section 4 (the non-stratified analysis of the rosiglitazione data set, using a weakly informative Cauchy(0.0, $\mathrm{Cauchy}(0.0,\,0.37)$ prior for the effect, and the analysis using a vague prior for the effect for the oncology data). In both cases, a $\mathrm{Halfnormal}(0.5)$ prior was specified for $\eta$.

### v2

@@ 6 Discussion
In evidence synthesis and meta-analysis contexts, the one-to-one correspondence of results from separate experiments or studies is often questionable, due to differences in designs, follow-up times, or definition of AEs. In addition, differences in capturing, adjudicating and reporting clinical events can increase between-study heterogeneity. The anticipation and consideration of (potential) heterogeneity between data sources hence is commonly advocated.44, 45 On the technical side, implementation of analysis models accounting for heterogeneity usually leads to more cautious or conservative inferences.34, 35 We have extended Holzhauer’s proposed meta-analysis model for synthesizing adverse event counts3 to incorporate a random-effect (RE) component in order to safeguard against overly optimistic or overconfident conclusions. When Holzhauer presented the original model, it was applied in an example case18 that had already triggered some controversy,26, 46, 47 and he was able to advance the discussion by leveraging additional external evidence via an informative MAP prior.4 We were able to demonstrate that the consideration of heterogeneity and use of an RE model also makes a difference here; using the rosiglitazone example as well as an additional application in oncology, we could show the impact of model specification details on the conclusions for these particular applications as well as in some limited Monte Carlo studies. Use of meta-analysis methods as well as consideration of heterogeneity between data sources allows to broaden the evidence base for decision making beyond the limits of single clinical trials that are usually powered with a focus on certain specific efficacy endpoints. Consideration of additional data will allow for more definite statements about harms and benefits, while allowing for exploratory investigations into wider ranges of questions. In the examples, very different disease areas were considered. Depending on the disease, only very specific AEs of interst qualify for such summaries as otherwise the noise from the anticipated events (i.e. those seen in the population of interest regardless of investigational treatment) may be too strong.
@@
Note that model specification in terms of a random “control” response (here: $\lambda_{i0}$) and a random treatment effect ($\varphi_{i}$) implies greater variability in the treatment responses compared to the control arms. In general, this is an often neglected feature, in the oncology example (Section 4.2), this was explicitly exploited in order to allow for greater variability for the (rather heterogeneous) control arms. Alternatively, one could specify a “symmetric” model variation as in Jackson et al. (2018),48 which may be preferrable in some applications. Another common issue in oncology applications is the use of active comparators, so that the extension to network meta-analysis methods may be an option; as long as only two-armed trials are included, this may be accommodated in a rather simple meta-regression model.49 Another aspect in which a more realistic model specification might be worth considering is a possible extension to a competing-risks model.50
The approach for adverse-event analysis considered here largely relies on the availability of relevant data and their reporting to a common standard. Unfortunately this is often not the case, hindering the assessment of a treatment’s risks and benefits. The analysis of clinical events and the comparison across trials would be much easier if the relevant data (such as HRs) were more commonly available; the establishment of consistent reporting standards would be very helpful in this context.51
The borrowing mechanism implemented in order to include external, historical data may be extended to potentially address additional requirements. Robustification was accomplished here via a mixture prior,4 whose design features (vague component, weights) need to be chosen carefully in order to avoid unintended behaviour. Alternatively, this may be addressed by explicitly implementing a more sophisticated joint likelihood reflecting the exact assumptions on the relationship between external and current data in terms of a bias allowance model.52, 53
In the simulations and example applications we used vague prior specifications (such as a variance of $100^{2}$ for location parameter priors). In the present context, these were meant to reflect conservative defaults, while in any specific practical application, there may be some prior information providing reasonable constraints that may be worth considering in the analysis.54 For location parameters, it may be most sensible to specify priors as “neutral” (e.g., with zero mean for log-HRs) and vague, where the magnitude of the prior variance may be motivated by the range of effect sizes considered plausible, or relative to the variance that would correspond to the information conveyed by a single observation.55 Variance (heterogeneity) parameters may be specified based on general considerations of plausible amounts of variation between studies,28 or sometimes also with reference to relevant empirical evidence.56, 57
When considering time-to-event data, an exponential model for waiting times may be an obvious starting point, while the validity of such an assumption may sometimes be questionable. In our simulations, we also considered scenarios with violations of this assumption, which are shown in Appendix C. As expected, operating characteristics deteriorate when the modelling assumption is not met; in practice, in case specific assumptions on the time-to-event model may be hard to justify, implementation of more flexible families (like Weibull, Lomax or piecewise exponential distributions) may be desirable. While in the applications presented here we used an implementation using Stan,39 analyses may also be formulated using the brms R package, which may also simplify extensions of the models considered here.58
@@ Appendix A Heterogeneity posteriors
Figure 8 illustrates the posteriors for the heterogeneity parameters ($\eta$) for two selected analyses of the rosiglitazone and oncology data sets from Section 4 (the non-stratified analysis of the rosiglitazione data set, using a weakly informative $\mathrm{Cauchy}(0.0,\,0.37)$ prior for the effect, and the analysis using a vague prior for the effect for the oncology data). In both cases, a $\mathrm{Halfnormal}(0.5)$ prior was specified for $\eta$.
Appendix B Point estimation
Table 7 illustrates the behaviour of point estimates (log-HRs, in terms of their means and standard deviations) for the simulations motivated by the rosiglitazone example (see Section 5.2.1.
Table 7: Performance of point estimates (in terms mean and standard deviations) of log-HRs for the simulation settings as described in Section 5.2.1.
simulation settings posterior mean log-HR: mean (st.dev.)
scenario log-HR $\varphi$ heterogeneity $\eta$ prior common effect (CE) random effects (RE)
1 0.00 0.00 (none) vague $0.007$ ($0.133$) 0.001-0.001 ($0.134$)
MAP $0.002$ ($0.090$) 0.006-0.006 ($0.092$)
2 0.00 0.10 (small) vague $0.002$ ($0.145$) 0.007-0.007 ($0.145$)
MAP 0.005-0.005 ($0.100$) 0.015-0.015 ($0.101$)
3 0.00 0.20 (moderate) vague $0.020$ ($0.159$) $0.003$ ($0.159$)
MAP $0.009$ ($0.126$) 0.008-0.008 ($0.124$)
4 0.00 0.40 (substantial) vague $0.065$ ($0.222$) $0.016$ ($0.207$)
MAP $0.056$ ($0.201$) $0.004$ ($0.185$)
5 0.00 0.80 (large) vague $0.198$ ($0.390$) $0.001$ ($0.345$)
MAP $0.193$ ($0.385$) $0.008$ ($0.334$)
6 0.25 0.00 (none) vague $0.256$ ($0.130$) $0.249$ ($0.131$)
MAP $0.246$ ($0.085$) $0.239$ ($0.086$)
7 0.25 0.10 (small) vague $0.250$ ($0.143$) $0.243$ ($0.144$)
MAP $0.247$ ($0.098$) $0.239$ ($0.099$)
8 0.25 0.20 (moderate) vague $0.268$ ($0.154$) $0.253$ ($0.152$)
MAP $0.262$ ($0.122$) $0.247$ ($0.120$)
9 0.25 0.40 (substantial) vague $0.310$ ($0.220$) $0.254$ ($0.207$)
MAP $0.298$ ($0.201$) $0.244$ ($0.186$)
10 0.25 0.80 (large) vague $0.441$ ($0.409$) $0.239$ ($0.364$)
MAP $0.435$ ($0.407$) $0.243$ ($0.351$)
Similarly, Table 8 eveluates point estimation in the simulation settings motivated by the oncology example (see Section 5.3.1).
Table 8: Performance of point estimates (in terms mean and standard deviations) of log-HRs for the simulation settings as described in Section 5.3.1.
simulation settings posterior mean log-HR: mean (st.dev.)
scenario log-HR $\varphi$ heterogeneity $\eta$ prior common effect (CE) random effects (RE)
1 0.00 0.00 (none) vague $0.011$ ($0.131$) $0.028$ ($0.138$)
2 0.00 0.10 (small) vague $0.006$ ($0.159$) $0.028$ ($0.164$)
3 0.00 0.20 (moderate) vague 0.019-0.019 ($0.167$) $0.005$ ($0.166$)
4 0.00 0.40 (substantial) vague 0.054-0.054 ($0.237$) 0.005-0.005 ($0.215$)
5 0.00 0.80 (large) vague 0.201-0.201 ($0.421$) 0.032-0.032 ($0.335$)
6 0.50 0.00 (none) vague $0.505$ ($0.158$) $0.531$ ($0.172$)
7 0.50 0.10 (small) vague $0.503$ ($0.172$) $0.529$ ($0.179$)
8 0.50 0.20 (moderate) vague $0.489$ ($0.180$) $0.516$ ($0.180$)
9 0.50 0.40 (substantial) vague $0.438$ ($0.245$) $0.491$ ($0.228$)
10 0.50 0.80 (large) vague $0.289$ ($0.427$) $0.460$ ($0.337$)
Appendix C Non-exponential time-to-event
In order to investigate departures from the assumption of exponentially distributed waiting times for events (equation (1)), we initiated limited simulations based on Weibull distributed times as variations of the simulations described in Sections 5.2.1 and 5.3.1. As expected, with violated assumptions the operating characteristics tend to worsen; details on coverage probabilities, type-I error and power are provieded in Tables 9 and 10. It should also be noted that in our implementation, use of the Weibull distribution for data generation led to increased rates of non-convergence for the MCMC, suggesting that longer burn-in and sampling times may be appropriate.
Table 9: Coverage probabilities, type-I error and power for variations of the simulation scenarios described in Section 5.2.1. Heterogeneity is at $\eta=0.2$ throughout, and Weibull-distributed times-to-event were simulated with several settings for the shape parameter.
simulation settings coverage (%) type-I error (%) power (%)
scenario log-HR $\varphi$ Weibull shape prior CE RE CE RE CE RE
3 0.00 1.5 vague 66.4 97.6 33.6 2.4
MAP 64.7 90.6 35.3 9.4
3 0.00 2.0 vague 64.9 97.6 35.1 2.4
MAP 61.2 89.4 38.8 10.6
3 0.00 2.5 vague 63.2 98.5 36.8 1.5
MAP 61.3 87.8 38.7 12.2
8 0.25 1.5 vague 57.2 96.8 68.1 10.4
MAP 56.5 93.6 66.0 13.6
8 0.25 2.0 vague 47.0 96.2 69.9 7.7
MAP 48.9 91.4 66.0 16.5
8 0.25 2.5 vague 40.8 97.4 72.4 6.6
MAP 42.8 86.1 68.8 23.2
Table 10: Coverage probabilities, type-I error and power for variations of the simulation setups described in Section 5.3.1. Heterogeneity is at $\eta=0.2$ throughout, and times-to-event are simulated based on a Weibull distribution with shape $1.5$.
simulation settings coverage (%) type-I error (%) power (%)
scenario log-HR $\varphi$ Weibull shape prior CE RE CE RE CE RE
3 0.0 1.5 vague 70.8 93.4 29.2 6.6
8 0.5 1.5 vague 67.9 85.1 66.1 55.6